QUESTION 3 (30 MARKS) Ahmad Best Sdn Bhd manufactures a popular ginseng canned drinks that are sold throughout Malaysia. The company sells an average of 1,200,000 cans per year. The following is the projected income statement for year 2018. Ahmad Best Sdn Bhd Projected Income Statement for the year 2018 Estimated cans of sales Selling price per can Estimated sales Estimated cans of sales Variable costs per can Estimated total variable costs Estimated fixed costs: 1,200,000 S 1.50 s 1,800,000 1,200,000 0.50 (S 600,000) Salaries Rent on building 86,000 64,000 65,800 47,000 Depreciation on equipment Other 54,200 Estimated total fixed cost Estimated profit per year S 317,000 $ 883,000 y expects volume of sales to decrease by 50% in 2018 due to new rivals in the markets. To reduce the elasticity of demand for its products, the company is considering two alternatives: I. To decrease each component of fixed costs by 20% but not to affect variable costs and sales price. 2. Not to affect fixed costs and variable costs but decrease selling price to $ 1.20 Required: Calculate the break-even point (in cans and S) of ginseng drinks, and the percentage of the firms margin of safety in 2018. (a) 9 marks) Calculate the break-even point in (cans and S) of ginseng drinks and the percentage of the fims margin of safety if the company chooses the first alternative in 2018 after considering the decrease in sales. (b) 9 marks) How many cans would have to be sold for the company in 2018 to earn a profit of $2,000,000 if the company chooses the second altermative? (c) (4 marks) ACCTS2IOMSept2018 Page 4 of 7
Answer #1

a. Breakeven point in cans = Fixed Cost / (Revenue per unit - variable cost per unit) = $317,000/($1.50-$0.50) = 317,000 cans.

Breakeven point in $ = Breakeven point in Units * Sales Price per unit = 317000 * 1.50 = $475,500.

Percentage of Margin of Safety of 2018 = (Actual Sales - Breakeven Sales) / Actual Sales * 100

= (1,800,000 - 475,500) / 1,800,000 * 100 = 73.58%

b. Company is planning to decrease each Fixed Cost component by 20%. So the revised total fixed Cost become 253,600 [317,000 * (1-0.20)].

Breakeven point in cans = Fixed Cost / (Revenue per unit - variable cost per unit) = $253,600/($1.50-$0.50) = 253,600 cans.

Breakeven point in $ = Breakeven point in Units * Sales Price per unit = 253,600 * 1.50 = $380,400.

Sales in projected to decrease by 50% due to new rivals in the market. So here the projected sales become 900,000 [1,800,000 * (1-0.50)]

Percentage of Margin of Safety of 2018 = (Actual Sales - Breakeven Sales) / Actual Sales * 100

= (900,000 - 380,400) / 900,000 * 100 = 57.73%

c. Here the company is decreasing selling price to 1.20$. So the contribution per unit becomes $0.70. The company wants to earn profit of $2,000,000 and the fixed cost is $317,000.

So the number of units the company will have to sell to earn $2,000,000 profit = ($2,000,000 + $317,000) / 0.70 = 3,310,000 cans.